Take a piece of string 50 cm long. Join the ends,'
how many different rectangles can you form so that the dimensions are whole numbers and the width is atleast 5cm.


please answer immediately
i really really need it, i hope after 45 min.there will be an answer.. :)
Thanks in advance!

Answer by Alan3354(69443)

2L + 2W = 50 (the perimeter)
L + W = 25
W = 5 to 24
5 to 24 inclucive is 20 integer values